# Matroid Products via Submodular Coupling

### v1

@@ 1 Introduction
One can think of the tensor product of two matroids as the “freest” among their possible products as it provides an unrestricted way to combine two matroids while preserving structure along both dimensions, meaning that the product of independent sets is also independent. However, since a tensor product does not exist for every pair of matroids, Mason [30] raised the problem of whether a freest product exists among the possible products. Las Vergnas [25] answered this question in the negative, and showed that the matroid obtained via Dilworth completion as in [27, 30] is the freest quasi-product satisfying $r(\{x,x^{\prime}\}\times\{y,y^{\prime}\})\leq 3$ for all $x,x^{\prime}\in S_{1}$ and $y,y^{\prime}\in S_{2}$, where $r$ denotes the rank function of the product matroid.
In summary, while the marginals of both couplings of probability measures and matroid tensor products match the original measures or rank functions up to a constant factor, they differ in key structural aspects. Couplings do not require measures to appear on sets of the form $\{x\}\times S_{2}$ or $S_{1}\times\{y\}$ called fibers, whereas matroid product constructions enforce this as a strict condition. Additionally, in probability couplings, the measure on the product space equals the product of the original measures, while for matroids, the rank of the product is typically lower than the product of the individual ranks.
@@ 1.1 Our results and techniques
Let $\varphi_{1}\colon 2^{S_{1}}\to\mathbb{R}$ and $\varphi_{2}\colon 2^{S_{2}}\to\mathbb{R}$ be set functions defined over finite ground sets $S_{1}$ and $S_{2}$, respectively. The notion of a tensor product can be naturally extended to set functions by calling a function $\varphi\colon 2^{S_{1}\times S_{2}}\to\mathbb{R}$ a tensor product of $\varphi_{1}$ and $\varphi_{2}$ if $\varphi(X_{1}\times X_{2})=\varphi_{1}(X_{1})\cdot\varphi_{2}(X_{2})$ holds for every $X_{1}\subseteq S_{1}$ and $X_{2}\subseteq S_{2}$. As a new concept, we call $\varphi$ a coupling of $\varphi_{1}$ and $\varphi_{2}$ if $\varphi(X_{1}\times S_{2})=\varphi_{1}(X_{1})\cdot\varphi_{2}(S_{2})$ for every $X_{1}\subseteq S_{1}$ and $\varphi(S_{1}\times X_{2})=\varphi_{1}(S_{1})\cdot\varphi_{2}(X_{2})$ for every $X_{2}\subseteq S_{2}$. In other words, a coupling is a set function on the product set, where the projections onto each coordinate return the corresponding $\varphi_{i}$, up to a constant multiplier.
@@
Let $\varphi$ be a set function over some finite ground set $S$, and let $\mathcal{Q}=(S_{1},\dots,S_{q})$ be a partition of $S$ into $q$ possibly empty parts. The quotient of $\varphi$ with respect to $\mathcal{Q}$ is a set function $\varphi_{\mathcal{Q}}$ over $S_{\mathcal{Q}}=\{s_{1},\dots,s_{q}\}$ defined by $\varphi_{\mathcal{Q}}(X)=\varphi(\bigcup_{s_{i}\in X}S_{i})$ for $X\subseteq S_{\mathcal{Q}}$. In particular, if $\varphi\colon 2^{S_{1}\times S_{2}}\to\mathbb{R}$ is a coupling of $\varphi_{1}\colon 2^{S_{1}}\to\mathbb{R}$ and $\varphi_{2}\colon 2^{S_{2}}\to\mathbb{R}$ and $\mathcal{Q}_{1}$ and $\mathcal{Q}_{2}$ are the partitions of $S_{1}\times S_{2}$ into fibers of the form $\{x\}\times S_{2}$ and $S_{1}\times\{y\}$, respectively, then $\varphi_{1}=\varphi_{2}(S_{2})\cdot\varphi_{\mathcal{Q}_{1}}$ and $\varphi_{2}=\varphi_{1}(S_{1})\cdot\varphi_{\mathcal{Q}_{2}}$. The positive results of Section 3 motivate the following question: Given a property of set functions, is there a function that is universal with respect to that property, in the sense that every function exhibiting the property can be obtained as its quotient?
@@ 2 Preliminaries
For basic definitions on matroids, we refer the reader to [32]. A matroid $M=(S,r)$ is defined by its finite ground set $S$ and its rank function $r\colon 2^{S}\to\mathbb{Z}_{+}$ that satisfies the rank axioms: (R1) $r(\emptyset)=0$, (R2) $X\subseteq Y\Rightarrow r(X)\leq r(Y)$, (R3) $r(X)\leq|X|$, and (R4) $r(X)+r(Y)\geq r(X\cap Y)+r(X\cup Y)$. Here, (R1) is a standard normalizing assumption, while (R2) requires the rank function to be increasing, (R3) its subcardinality, and (R4) its submodularity. The rank of the matroid is $r(S)$, and by the normalized rank function of $M$ we mean the set function $r/r(S)$. A subset $X\subseteq S$ is called independent if $|X|=r(X)$. An inclusionwise minimal non-independent set forms a circuit, while a loop is a circuit consisting of a single element. A flat is a set $F\subseteq S$ such that $r(F+e)=r(F)$ holds for every $e\in S\setminus F$. A matroid is called modular if $r(F_{1})+r(F_{2})=r(F_{1}\cap F_{2})+r(F_{1}\cup F_{2})$ for any pair of flats $F_{1},F_{2}$. For sets $X,Y\subseteq S$, we say that $X$ spans $Y$ in $M$ if $r(X\cup Y)=r(X)$.
@@
The following is a fundamental result in polyhedral combinatorics, see e.g. [17, Corollary 14.2.3].
@@
It is not difficult to see that $\varphi_{A}$ is a normalized coverage function. Choquet [9, Section 43] showed that these set functions correspond to the extreme rays of the convex cone of coverage functions; see also [29] for a different proof.
@@ 3.3.1 Matroid couplings and amalgams
$r(Z)=\min_{W\supseteq Z}\left\{\sum_{e_{1}\in B_{1}}r_{2}(\pi_{2}(W_{e_{1}}))+\sum_{e_{2}\in B_{2}}r_{1}(\pi_{1}(W^{e_{2}})-|W\cap(B_{1}\times B_{2})|\right\}.$
@@
There exists a matroid $M$ over $[n]^{N_{r,n}}$ of rank $r^{N_{r,n}}$ such that any rank-$r$ matroid on $n$ is a projection of $M$ to one of its coordinates.
@@
Let $N_{0}=(T_{1}\cap T_{2},r_{0}^{\prime})$, $N_{1}=(T_{1},r^{\prime}_{1})$ and $N_{2}=(T_{2},r^{\prime}_{2})$ be matroids such that $N_{1}|(T_{1}\cap T_{2})=N_{2}|(T_{1}\cap T_{2})=N_{0}$, and let $T\coloneqq T_{1}\cup T_{2}$. A matroid $N=(T,r)$ is called an amalgam of $N_{1}$ and $N_{2}$ if $N|T_{1}=N_{1}$ and $N|T_{2}=N_{2}$. An amalgam $N$ is called free if for any amalgam $N^{\prime}$, each independent set of $N^{\prime}$ is independent in $N$. The following result appears, for example, in [32, Proposition 11.4.2].
@@
Let $N_{0}=(T_{1}\cap T_{2},r^{\prime})$, $N_{1}=(T_{1},r^{\prime}_{2})$, and $N_{2}=(T_{2},r^{\prime}_{1})$ be matroids such that $N_{1}|(T_{1}\cap T_{2})=N_{2}|(T_{1}\cap T_{2})=N_{0}$, and let $T\coloneqq T_{1}\cup T_{2}$. Define $r\colon 2^{T}\to\mathbb{Z}_{+}$ by
@@ 3.3.3 Linear matroids and Ingleton’s inequality
For any fixed ground set $S_{1}$ and fixed matroid $N=(S_{2},r_{2})$, the set of submodular (or polymatroid) functions $\varphi_{1}\colon 2^{S_{1}}\to\mathbb{R}$ which have a tensor product with $r_{2}$ forms a polyhedron in $\mathbb{R}^{2^{S_{1}}}$, that is, they can be described as functions $\varphi\colon 2^{S_{1}}\to\mathbb{R}$ satisfying certain linear inequalities. If we choose $N$ to be a regular matroid, then each of these inequalities is a linear rank inequality, that is, a linear inequality satisfied by the rank function of any representable matroid. Ingleton’s inequality (7) was the first known linear rank inequality, and since then infinite families of independent linear rank inequalities has been found [11, 23]. Interestingly, all currently known linear rank inequalities follow from the so-called common information property of the rank function of representable matroids [11, 4], though other linear inequalities have been found that are satisfied by rank functions of matroids representable over fields of certain characteristics but not by all representable matroids [12]. For defining the common information property, let us call every polymatroid function 00-CI-compliant, and call a polymatroid function $f\colon 2^{S}\to\mathbb{R}$ is $k$-CI-compliant if for any subsets $A,B\subseteq S$, there exists a $(k-1)$-compliant function $f^{\prime}\colon 2^{S^{\prime}}\to\mathbb{R}$ and a subset $Z\subseteq S^{\prime}$ such that $S\subseteq S^{\prime}$, $f^{\prime}(X)=f(X)$ for all $X\subseteq S$, $f^{\prime}(A\cup Z)=f(A)$, $f^{\prime}(B\cup Z)=f(B)$, and $f^{\prime}(Z)=f(A)+f(B)-f(A\cup B)$. Then, a polymatroid satisfies the common information property if it is $k$-CI-compliant for each integer $k\geq 0$. Note that the common information property is not sufficient for the representability of a matroid, e.g. all rank-3 matroids satisfy it [4] while e.g. the non-Desargues matroid is not representable [21]. It would be interesting to see the relations between the class of polymatroid functions having tensor products with $U_{2,3}$, the class of polymatroid functions having tensor product with all regular matroids, and the class of polymatroid functions satisfying the common information property.
@@ 3.4 Coverage and $k$-alternating functions
In [6, Example 4.17], it is shown that the rank function of a matroid is a coverage function if and only if it is a partition matroid. Combining this observation with Theorem 3.16 implies that any matroid has a tensor product with any partition matroid, which is a well-known result on tensor products.
@@ 3.5 Extension to the infinite case
for all $Z\in\mathcal{B}$. While this definition does not look symmetric at first, note that $\hat{\alpha}_{1}(\alpha_{2}(\pi_{2}(Z_{x})))$ and $\hat{\alpha}_{2}(\alpha_{1}(\pi_{1}(Z^{y})))$ agree on the semi-algebra of product sets and thus agree on the generated algebra, see e.g. [33, Theorem 3.5.1 (ii)]. Observe that if $Z=Y_{1}\times Y_{2}$ for some $Y_{1}\in\mathcal{B}_{1}$ and $Y_{2}\in\mathcal{B}_{2}$, then
@@
if $Y_{1}=J_{1}$ or $Y_{2}=J_{2}$, hence $b$ is indeed a coupling. To prove submodularity, note that for fix $x\in J_{1}$ and $y\in J_{2}$, the functions $\varphi_{2}(\pi_{2}(Z_{x}))$ and $\varphi_{1}(\pi_{1}(Z^{y}))$ are submodular while $\alpha_{2}(\pi_{2}(Z_{x}))$ is modular, hence $b$ is submodular by the linearity and positivity of $\alpha_{1}$ and $\alpha_{2}$. ∎
To extend the result of 3.2, we first have to define the measurable analogue of a base polytope. Let $(J,\mathcal{B})$ be a set algebra and $\varphi\colon\mathcal{B}\to\mathbb{R}_{+}$ be a polymatroid function. We say that a nonnegative charge $\alpha\colon\mathcal{B}\to\mathbb{R}_{+}$ is a minorizing charge if $\alpha(A)\leq\varphi(A)$ for all $A\in\mathcal{B}$. A minorizing charge is called basic if $\alpha(J)=\varphi(J)$. The set of all basic minorizing charges is denoted by $\operatorname{bmm}(\varphi)$. It was proved in [29, Corollary 5.3.] that $\operatorname{bmm}(\varphi)$ is nonempty.
@@
We have $b(\emptyset)=0$ and $b(Y_{1}\times Y_{2})\leq\varphi_{1}(Y_{1})\cdot\varphi_{2}(Y_{2})$ for each $Y_{1}\in\mathcal{B}_{1}$ and $Y_{2}\in\mathcal{B}_{2}$, where equality holds if $\varphi_{1}(Y_{1})=\mu_{1}(Y_{1})$ or $\varphi_{2}(Y_{2})=\mu_{2}(Y_{2})$.
@@
The equality $b(\emptyset)=0$ follows from $\varphi_{1}(\emptyset)=\varphi_{2}(\emptyset)=0$. For each $Y_{1}\in\mathcal{B}_{1}$ and $Y_{2}\in\mathcal{B}_{2}$, $\mu_{1}(Y_{1})\leq\varphi_{1}(Y_{1})$ and $\mu_{2}(Y_{2})\leq\varphi_{2}(Y_{2})$, thus we get
@@
We need to show that $b(Z)\geq b(Y_{1}\times Y_{2})$ holds if $Y_{1}\times Y_{2}\subseteq Z\in\mathcal{B}$. Using that $\varphi_{2}$ is increasing and $\mu_{2}(X)\leq\varphi_{2}(X)$ holds for any $X\in\mathcal{B}_{2}$, we get

### v2

@@ Abstract
 [^1] An extended abstract [6] of the paper appeared at 57th Annual ACM Symposium on Theory of Computing (STOC 2025).
@@ 1 Introduction
One can think of the tensor product of two matroids as a construction where the product of independent sets is itself independent. However, since a tensor product does not exist for every pair of matroids, Mason [31] raised the question of whether a freest product exists among the possible products. Las Vergnas [26] answered this question in the negative and showed that the matroid obtained via Dilworth truncation as in [28, 31] is the freest quasi-product among those satisfying $r(\{x,x^{\prime}\}\times\{y,y^{\prime}\})\leq 3$ for all $x,x^{\prime}\in S_{1}$ and $y,y^{\prime}\in S_{2}$, where $r$ denotes the rank function of the product matroid.
@@ 1.1 Our results and techniques
Let $\varphi_{1}\colon 2^{S_{1}}\to\mathbb{R}$ and $\varphi_{2}\colon 2^{S_{2}}\to\mathbb{R}$ be set functions defined over finite ground sets $S_{1}$ and $S_{2}$, respectively. The notion of a tensor product can be naturally extended to set functions by calling a function $\varphi\colon 2^{S_{1}\times S_{2}}\to\mathbb{R}$ a tensor product of $\varphi_{1}$ and $\varphi_{2}$ if $\varphi(X_{1}\times X_{2})=\varphi_{1}(X_{1})\cdot\varphi_{2}(X_{2})$ holds for every $X_{1}\subseteq S_{1}$ and $X_{2}\subseteq S_{2}$. As a new concept, we call $\varphi$ a coupling of $\varphi_{1}$ and $\varphi_{2}$ if $\varphi(X_{1}\times S_{2})=\varphi_{1}(X_{1})\cdot\varphi_{2}(S_{2})$ for every $X_{1}\subseteq S_{1}$ and $\varphi(S_{1}\times X_{2})=\varphi_{1}(S_{1})\cdot\varphi_{2}(X_{2})$ for every $X_{2}\subseteq S_{2}$. In other words, a coupling is a set function on the product set, where the projections onto each coordinate return the corresponding $\varphi_{i}$, up to a constant multiplier. It is worth emphasizing that neither a tensor product nor a coupling is uniquely determined from the factors $\varphi_{1}$ and $\varphi_{2}$.
@@
Let $\varphi$ be a set function over some finite ground set $S$, and let $\mathcal{Q}=(S_{1},\dots,S_{q})$ be a partition of $S$ into $q$ possibly empty parts. The quotient of $\varphi$ with respect to $\mathcal{Q}$ is a set function $\varphi_{\mathcal{Q}}$ over $S_{\mathcal{Q}}=\{s_{1},\dots,s_{q}\}$ defined by $\varphi_{\mathcal{Q}}(X)=\varphi(\bigcup_{s_{i}\in X}S_{i})$ for $X\subseteq S_{\mathcal{Q}}$. In particular, if $\varphi\colon 2^{S_{1}\times S_{2}}\to\mathbb{R}$ is a coupling of $\varphi_{1}\colon 2^{S_{1}}\to\mathbb{R}$ and $\varphi_{2}\colon 2^{S_{2}}\to\mathbb{R}$ and $\mathcal{Q}_{1}$ and $\mathcal{Q}_{2}$ are the partitions of $S_{1}\times S_{2}$ into fibers of the form $\{x\}\times S_{2}$ and $S_{1}\times\{y\}$, respectively, then $\varphi_{1}=\varphi_{\mathcal{Q}_{1}}/\varphi_{2}(S_{2})$ and $\varphi_{2}=\varphi_{\mathcal{Q}_{2}}/\varphi_{1}(S_{1})$. The positive results of Section 3 motivate the following question: Given a property of set functions, is there a function that is universal with respect to that property, in the sense that every function exhibiting the property can be obtained as its quotient?
@@ 2 Preliminaries
For basic definitions on matroids, we refer the reader to [33]. A matroid $M=(S,r)$ is defined by its finite ground set $S$ and its rank function $r\colon 2^{S}\to\mathbb{Z}_{+}$ that satisfies the rank axioms: (R1) $r(\emptyset)=0$, (R2) $X\subseteq Y\Rightarrow r(X)\leq r(Y)$, (R3) $r(X)\leq|X|$, and (R4) $r(X)+r(Y)\geq r(X\cap Y)+r(X\cup Y)$. Here, (R1) is a standard normalizing assumption, while (R2) requires the rank function to be increasing, (R3) its subcardinality, and (R4) its submodularity. The rank of the matroid is $r(S)$, and by the normalized rank function of $M$ we mean the set function $r/r(S)$. A subset $X\subseteq S$ is called independent if $|X|=r(X)$. An inclusionwise minimal non-independent set forms a circuit, while a loop is a circuit consisting of a single element. A flat is a set $F\subseteq S$ such that $r(F+e)=r(F)+1$ holds for every $e\in S\setminus F$. A matroid is called modular if $r(F_{1})+r(F_{2})=r(F_{1}\cap F_{2})+r(F_{1}\cup F_{2})$ for any pair of flats $F_{1},F_{2}$. For sets $X,Y\subseteq S$, we say that $X$ spans $Y$ in $M$ if $r(X\cup Y)=r(X)$.
@@
The following is a fundamental result in polyhedral combinatorics, see e.g. [18, Corollary 14.2.3].
@@
It is not difficult to see that $\varphi_{A}$ is a normalized coverage function. Choquet [10, Section 43] showed that these set functions correspond to the extreme rays of the convex cone of coverage functions; see also [30] for a different proof.
@@ 3.3.1 Matroid couplings and amalgams
$r(Z)=\min_{W\supseteq Z}\left\{\sum_{e_{1}\in B_{1}}r_{2}(\pi_{2}(W_{e_{1}}))+\sum_{e_{2}\in B_{2}}r_{1}(\pi_{1}(W^{e_{2}}))-|W\cap(B_{1}\times B_{2})|\right\}.$
@@
There exists a matroid $M$ over $[n]^{N_{r,n}}$ of rank $r^{N_{r,n}}$ such that any rank-$r$ matroid on $[n]$ is a projection of $M$ to one of its coordinates.
@@
Let $N_{0}=(T_{1}\cap T_{2},r_{0}^{\prime})$, $N_{1}=(T_{1},r^{\prime}_{1})$ and $N_{2}=(T_{2},r^{\prime}_{2})$ be matroids such that $N_{1}|(T_{1}\cap T_{2})=N_{2}|(T_{1}\cap T_{2})=N_{0}$, and let $T\coloneqq T_{1}\cup T_{2}$. A matroid $N=(T,r)$ is called an amalgam of $N_{1}$ and $N_{2}$ if $N|T_{1}=N_{1}$ and $N|T_{2}=N_{2}$. An amalgam $N$ is called free if for any amalgam $N^{\prime}$, each independent set of $N^{\prime}$ is independent in $N$. The following result appears, for example, in [33, Proposition 11.4.2].
@@
Let $N_{0}=(T_{1}\cap T_{2},r^{\prime})$, $N_{1}=(T_{1},r^{\prime}_{1})$, and $N_{2}=(T_{2},r^{\prime}_{2})$ be matroids such that $N_{1}|(T_{1}\cap T_{2})=N_{2}|(T_{1}\cap T_{2})=N_{0}$, and let $T\coloneqq T_{1}\cup T_{2}$. Define $r\colon 2^{T}\to\mathbb{Z}_{+}$ by
@@ 3.3.3 Linear matroids and Ingleton’s inequality
For any fixed ground set $S_{1}$ and fixed matroid $N=(S_{2},r_{2})$, the set of submodular (or polymatroid) functions $\varphi_{1}\colon 2^{S_{1}}\to\mathbb{R}$ which have a tensor product with $r_{2}$ forms a polyhedron in $\mathbb{R}^{2^{S_{1}}}$, that is, they can be described as functions $\varphi_{1}\colon 2^{S_{1}}\to\mathbb{R}$ satisfying certain linear inequalities. If we choose $N$ to be a regular matroid, then each of these inequalities is a linear rank inequality, that is, a linear inequality satisfied by the rank function of any representable matroid. Ingleton’s inequality (7) was the first known linear rank inequality, and since then infinite families of independent linear rank inequalities has been found [12, 24]. Interestingly, all currently known linear rank inequalities follow from the so-called common information property of the rank function of representable matroids [12, 4], though other linear inequalities have been found that are satisfied by rank functions of matroids representable over fields of certain characteristics but not by all representable matroids [13]. For defining the common information property, let us call every polymatroid function 00-CI-compliant, and call a polymatroid function $f\colon 2^{S}\to\mathbb{R}$ is $k$-CI-compliant if for any subsets $A,B\subseteq S$, there exists a $(k-1)$-compliant function $f^{\prime}\colon 2^{S^{\prime}}\to\mathbb{R}$ and a subset $Z\subseteq S^{\prime}$ such that $S\subseteq S^{\prime}$, $f^{\prime}(X)=f(X)$ for all $X\subseteq S$, $f^{\prime}(A\cup Z)=f(A)$, $f^{\prime}(B\cup Z)=f(B)$, and $f^{\prime}(Z)=f(A)+f(B)-f(A\cup B)$. Then, a polymatroid satisfies the common information property if it is $k$-CI-compliant for each integer $k\geq 0$. Note that the common information property is not sufficient for the representability of a matroid, e.g. all rank-3 matroids satisfy it [4] while e.g. the non-Desargues matroid is not representable [22]. It would be interesting to see the relations between the class of polymatroid functions having tensor products with $U_{2,3}$, the class of polymatroid functions having tensor product with all regular matroids, and the class of polymatroid functions satisfying the common information property.
@@ 3.4 Coverage and $k$-alternating functions
In [7, Example 4.17], it is shown that the rank function of a matroid is a coverage function if and only if it is a partition matroid. Combining this observation with Theorem 3.16 implies that any matroid has a tensor product with any partition matroid, which is a well-known result on tensor products.
@@ 3.5 Extension to the infinite case
for all $Z\in\mathcal{B}$. While this definition does not look symmetric at first, note that $\hat{\alpha}_{1}(\alpha_{2}(\pi_{2}(Z_{x})))$ and $\hat{\alpha}_{2}(\alpha_{1}(\pi_{1}(Z^{y})))$ agree on the semi-algebra of product sets and thus agree on the generated algebra, see e.g. [34, Theorem 3.5.1 (ii)]. Observe that if $Z=Y_{1}\times Y_{2}$ for some $Y_{1}\in\mathcal{B}_{1}$ and $Y_{2}\in\mathcal{B}_{2}$, then
@@
if $Y_{1}=J_{1}$ or $Y_{2}=J_{2}$, hence $b$ is indeed a coupling. To prove submodularity, note that for fix $x\in J_{1}$ and $y\in J_{2}$, the functions $\varphi_{2}(\pi_{2}(Z_{x}))$ and $\varphi_{1}(\pi_{1}(Z^{y}))$ are submodular while $\alpha_{2}(\pi_{2}(Z_{x}))$ is modular, hence $b$ is submodular by the linearity and nonnegativity of $\alpha_{1}$ and $\alpha_{2}$. ∎
To extend the result of 3.2, we first have to define the measurable analogue of a base polytope. Let $(J,\mathcal{B})$ be a set algebra and $\varphi\colon\mathcal{B}\to\mathbb{R}_{+}$ be a polymatroid function. We say that a nonnegative charge $\alpha\colon\mathcal{B}\to\mathbb{R}_{+}$ is a minorizing charge if $\alpha(A)\leq\varphi(A)$ for all $A\in\mathcal{B}$. A minorizing charge is called basic if $\alpha(J)=\varphi(J)$. The set of all basic minorizing charges is denoted by $\operatorname{bmm}(\varphi)$. It was proved in [30, Corollary 5.3.] that $\operatorname{bmm}(\varphi)$ is nonempty.
@@
We have $b(\emptyset)=0$ and $b(Y_{1}\times Y_{2})\leq\varphi_{1}(Y_{1})\cdot\varphi_{2}(Y_{2})$ for each $Y_{1}\in\mathcal{B}_{1}$ and $Y_{2}\in\mathcal{B}_{2}$, where equality holds if $\varphi_{1}(Y_{1})=\alpha_{1}(Y_{1})$ or $\varphi_{2}(Y_{2})=\alpha_{2}(Y_{2})$.
@@
The equality $b(\emptyset)=0$ follows from $\varphi_{1}(\emptyset)=\varphi_{2}(\emptyset)=0$. For each $Y_{1}\in\mathcal{B}_{1}$ and $Y_{2}\in\mathcal{B}_{2}$, $\alpha_{1}(Y_{1})\leq\varphi_{1}(Y_{1})$ and $\alpha_{2}(Y_{2})\leq\varphi_{2}(Y_{2})$, thus we get
$b(Y_{1}\times Y_{2})=\varphi_{1}(Y_{1})\cdot\varphi_{2}(Y_{2})-(\varphi_{1}(Y_{1})-\alpha_{1}(Y_{1}))\cdot(\varphi_{2}(Y_{2})-\alpha_{2}(Y_{2}))\leq\varphi_{1}(Y_{1})\cdot\varphi_{2}(Y_{2}).$
Here equality holds if $\varphi_{1}(Y_{1})=\alpha_{1}(Y_{1})$ or $\varphi_{2}(Y_{2})=\alpha_{2}(Y_{2})$. ∎
@@
We need to show that $b(Z)\geq b(Y_{1}\times Y_{2})$ holds if $Y_{1}\times Y_{2}\subseteq Z\in\mathcal{B}$. Using that $\varphi_{2}$ is increasing and $\alpha_{2}(X)\leq\varphi_{2}(X)$ holds for any $X\in\mathcal{B}_{2}$, we get